# Uni-X: Mitigating Modality Conflict with a Two-End-Separated Architecture for Unified Multimodal Models

### v2

@@ 5 Conclusions
We are committed to ensuring the reproducibility of our research. To this end, we provide comprehensive details throughout the paper and in the appendix.
Code. The source code for our Uni-X model architecture, training, and evaluation is made available at the following anonymous repository: https://anonymous.4open.science/r/Uni-X-Code-E5CD.
Architecture and Implementation. The detailed architecture of Uni-X is described in Section 3.2. Implementation details, including the base models used (Qwen2.5-1.5B and Qwen2.5-3B), the VQGAN tokenizer, and software dependencies are provided in Section 4.1. Details on our baseline implementations are available in Appendix A.2.
Datasets and Evaluation. All datasets used for pre-training and supervised fine-tuning are listed in Section 4.1. The evaluation benchmarks for understanding and generation tasks are also detailed in the same section.
Hyperparameters. Key hyperparameters for our main ablation study, including learning rate, batch size, and scheduler details, are specified in Section 4.3 to ensure a fair comparison.
We believe that the combination of our provided code, detailed architectural descriptions, dataset lists, and specific hyperparameters will enable the community to replicate our findings and build upon our work.
@@ Appendix A Appendix
A.1 Prompts of Image Generation
@@
A.2 Baseline Implementation Details
To ensure fair comparisons, we adapt baseline methods to the VQ+AR setting used in our study. For Mixture-of-Transformers (MoT) (Deng et al., 2025; Shi et al., 2025; Liao et al., 2025), the duplicated transformer is originally designed for image generation through diffusion. To remove the influence of diffusion and isolate architectural effects, we reconfigure the duplicated transformer to operate directly on image tokens. In this setup, the qkv sequences from the two transformers are concatenated within the attention module, allowing the model to incorporate visual information for both understanding and generation tasks. As a result, the MoT results reported in this paper reflect its effectiveness strictly within the VQ+AR paradigm, eliminating confounding factors introduced by diffusion-based processes.
A.3 More Evaluation Results on Image Generation Benchmark
@@
A.4 Gradient Conflict Analysis
@@
A.5 Inference Efficiency
We have conducted a comprehensive evaluation of inference efficiency on an H800 PCIe (350W) GPU. As shown in Table 10, Uni-X demonstrates superior throughput compared to standard autoregressive baselines.
Uni-X achieves high throughput (910.2 tokens/s) even compared to the original Qwen2.5-3B (975.2 tokens/s), despite the architectural changes and a higher number of parameters (4.5B vs 3B). This efficiency gain stems from the computational complexity of the attention mechanism in the separated layers.
@@ A.5 Inference Efficiency
Model # Params. Throughput $\uparrow$
Tokens/s Images/min
@@
A.6 Ablation Study on Ratio between Text and Vision.
@@
A.7 Use of Large Language Models

### v3

@@ 5 Conclusions
Acknowledgments
We would like to thank the anonymous reviewers for their insightful feedback and constructive suggestions, which have significantly improved the quality of this paper. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 62125201, U24B20174, and U25B6003.
@@ Acknowledgments
We are committed to ensuring the full reproducibility of our work. To facilitate verification and future extensions, we provide detailed descriptions of the model architecture, training setup, datasets, evaluation protocols, and implementation details throughout the main paper and appendix.
Code.
The complete source code for the Uni-X architecture, training pipeline, and evaluation scripts is publicly available at: https://github.com/CURRENTF/Uni-X. The repository includes model definitions, data preprocessing utilities, configuration files, and instructions for reproducing both ablation and scaling experiments.
Architecture and Implementation.
The description of the Uni-X architecture is provided in Section 3.2. Implementation details, including the base models (Qwen2.5-1.5B and Qwen2.5-3B), the VQGAN tokenizer configuration, training framework (Liquid and HuggingFace Transformers), optimization tools (FlashAttention 2 and DeepSpeed ZeRO2), and image generation settings (e.g., CFG = 4.0), are documented in Section 4.1. Adaptations made to baseline methods under the VQ+AR setting are fully described in Appendix A.3 to ensure fair and transparent comparison.
Datasets and Evaluation Benchmarks.
All pre-training datasets (including text and multimodal corpora) and supervised fine-tuning datasets are listed in Section 4.1, along with token statistics. Evaluation benchmarks for text-only tasks, image generation, multimodal understanding, and image editing are also described in Section 4.1 and corresponding result tables. Where prompt rewriting or evaluation-specific adjustments were applied (e.g., GenEval prompt expansion), these procedures are explicitly documented.
Hyperparameters and Training Settings.
Key training hyperparameters for ablation experiments, including learning rate, warmup ratio, batch size, and scheduler details, are specified in Section 4.1. Scaling experiments and total token counts (e.g., 140B / 240B settings) are also clearly reported in Section 4.2. These details ensure that experimental comparisons are conducted under fully specified and reproducible conditions.
We believe that the combination of open-source code, precise architectural definitions, transparent dataset descriptions, and explicitly reported training hyperparameters enables the research community to faithfully reproduce our results and build upon the Uni-X framework for future unified multimodal modeling research.
@@ Appendix A Appendix
A.1 Other Related Work
Other UMMs
Recent advancements in Unified Multimodal Models (UMMs) have explored various architectural and methodological innovations. For instance, OneCAT (Li et al., 2025b) proposes a decoder-only autoregressive architecture that utilizes modality-specific Mixture-of-Experts (MoE) structures to achieve efficient multimodal fusion without visual encoders. VARGPT (Zhuang et al., 2025) unifies visual understanding and generation through a seamless combination of next-token and next-scale prediction paradigms. To harmonize multi-granularity representations, the Harmon framework (Wu et al., 2025d) leverages a shared masked autoregressive encoder. Furthermore, MetaQueries (Pan et al., 2025) introduces learnable queries to transfer reasoning capabilities from frozen MLLMs to diffusion decoders.
UMMs for Efficient Agent
The capabilities of vision-language models are continuously evolving, driving their application in specialized vertical domains and complex agentic tasks. For example, the emerging “Deep Research” paradigm (Shi et al., 2025b), iterative self-incentivization frameworks (Shi et al., 2025c), and domain-specific benchmarks like DeepShop (Lyu et al., 2025b) highlight the critical importance of tool utilization and autonomous multi-step interactions in real-world scenarios. Currently, translating these capabilities into complex tasks often relies on cumbersome multi-agent or multi-model pipelines. Instances of this include cooperative multi-agent systems for entity recognition (Wang et al., 2025) and alignment optimization (Lyu et al., 2025a), as well as specialized fragmented models for fine-grained document grounding (Zhou et al., 2025), search and recommendation (Zhao et al., 2026), and preference optimization (Wu et al., 2025b). However, these multi-model pipelines inevitably generate extremely long contexts and heavy computational burdens. While these efficiency issues can be partially alleviated by inference acceleration and efficient training methods like OmniKV (Hao et al., 2025b), SnapKV (Li et al., 2024b), and Low-Rank Clone (Hao et al., 2025a), relying on fragmented systems lacks architectural elegance (Shen et al., 2026; Hao et al., 2026). In contrast, UMMs offer a more concise and efficient alternative, enabling a single unified model to seamlessly handle diverse reasoning, tool-calling, and multimodal tasks without the cumbersome overhead of complex multi-agent interactions.
A.2 Prompts of Image Generation
@@
A.3 Baseline Implementation Details
To ensure fair comparisons, we adapt baseline methods to the VQ+AR setting used in our study. For Mixture-of-Transformers (MoT) (Deng et al., 2025; Shi et al., 2025a; Liao et al., 2025), the duplicated transformer is originally designed for image generation through diffusion. To remove the influence of diffusion and isolate architectural effects, we reconfigure the duplicated transformer to operate directly on image tokens. In this setup, the qkv sequences from the two transformers are concatenated within the attention module, allowing the model to incorporate visual information for both understanding and generation tasks. As a result, the MoT results reported in this paper reflect its effectiveness strictly within the VQ+AR paradigm, eliminating confounding factors introduced by diffusion-based processes.
A.4 More Evaluation Results on Image Generation Benchmark
@@
A.5 Gradient Conflict Analysis
@@
A.6 Inference Efficiency
We have conducted a comprehensive evaluation of inference efficiency on an H800 PCIe (350W) GPU. As shown in Table 10, Uni-X demonstrates superior throughput compared to standard autoregressive baselines. Uni-X achieves high throughput (910.2 tokens/s) even compared to the original Qwen2.5-3B (975.2 tokens/s), despite the architectural changes and a higher number of parameters (4.5B vs 3B). This efficiency gain stems from the computational complexity of the attention mechanism in the separated layers.
@@ A.6 Inference Efficiency
Model # Params. Tokens/s Images/min
@@
A.7 Ablation Study on Ratio between Text and Vision.
@@
A.8 Use of Large Language Models